OpenStudy (anonymous):

Initally, a 100-liter tank contains a salt solution with concentration 0.5 kg/liter. A fresher solution with concentration 0.1 kg/liter flows into the tank at the rate of 4 liters/min. The contents of the tank are kept well-stirred, and the mixture flows out at the same rate it flows in. For the questions below, we measure t in minutes. Determine the concentration (in kg/L) in the tank as a function of t time .

OpenStudy (anonymous):

x' = flow in - flow out x'/100 = concentration at t I figured that out minutes after asking :p

OpenStudy (dumbcow):

this took me awhile but i think i got it C(t) = (50+0.4t)/(100+4t) this came from the salt coming in was 0.1*4t salt going out was C(t)*4t net salt = C(t)*4t - 0.1*4t initial salt content = 50 C(t) = (50 - net salt)/100